Alice and Bob get married and become the monopoly producer of
Drink Me™ in Wonderland. Market demand for Drink Me™ is given
by p = 40 − 2q and cost of production is C(q) = q
2
. Please calculate
monopoly price and quantity?
1
Expert's answer
2017-05-03T09:04:08-0400
If p = 40 − 2q and C(q) = q2, then the monopoly quantity can be calculated at MR = MC, so:
MR = TR' = (p*q)' = 40 - 4q,
MC = C' = 2q,
40 - 4q = 2q,
6q = 40,
q = 6.67 units.
The monopoly price is p = 40 - 2*6.67 = $26.66.